# Can Open-source LLMs Enhance Data Synthesis for Toxic Detection?: An Experimental Study

### v2

@@ 4.3 Error Analysis
We conducted an error analysis on the fine-tuned models to better understand the nature of their mistakes and identify areas for improvement. This analysis focuses on two key aspects: (1) generation errors and (2) classification errors.
@@
The data generation process was evaluated for repetition rates and the nature of outputs across different epochs and model versions for the hate and sex datasets. For example Mixtral_Mixture3 models were effective in generating targeted harmful data but occasionally produced random string like " You are 32\$\$2\$(2djs…" , which could skew downstream training. Additionally, repetitive outputs across epochs indicated potential overfitting in certain scenarios, suggesting the need for better strategies to ensure diversity and control the tone of generated content.
@@
We examined classification errors in the downstream MLP classifier trained on augmented data. Examples of missed classifications include statements like, "You are a disabled person," which might have been overlooked due to its less explicit tone compared to direct threats. Similarly, a statement like, "I’m sorry you’re deaf," may have been misclassified as non-harmful due to its seemingly apologetic phrasing. These cases highlight the difficulty of identifying subtle harmful content and underscore the need for classifiers capable of handling nuanced language.
@@ 5 Discussion
The results of this study highlight the complexities and challenges associated with generating and detecting harmful content using LLMs. Through a two-stage evaluation process, we examined the performance of various open-source LLMs and assessed the impact of fine-tuning on mixed datasets. Several key insights emerged: One of the primary challenges observed was the subtlety of harmful content. The fine-tuned models often failed to detect subtle harmful statements that were not overtly aggressive or violent. This limitation underscores the need for more sophisticated detection mechanisms capable of understanding nuanced language and contextual subtleties. Repetition and overfitting were also significant issues, particularly in the later epochs of the fine-tuned models. High repetition rates suggest that the models may be memorizing specific phrases rather than generalizing effectively across different harmful contexts. This overfitting reduces the diversity and overall effectiveness of harmful content generation. The data mixing strategy, which combined datasets of hate, violence, and sex with GPT-generated positive samples, showed promising results in improving both precision and recall. However, the models continued to struggle with maintaining a balance between specificity and generalization, especially when dealing with mixed types of harmful content.
6 Online Deployment
The models developed in this study have been successfully deployed in a real-world application by a leading cloud provider, offering valuable insights and automations for a range of tasks. Specifically, these models are used to enhance content moderation workflows, enabling faster and more accurate detection of harmful content in real-time scenarios. Their deployment has demonstrated the practicality of fine-tuned open-source LLMs in addressing challenges such as latency, scalability, and adaptability to diverse content types.
7 Conclusion
This study showcases the potential of open-source LLMs for harmful data synthesis and their practical applications in toxic content detection systems. By evaluating prompt engineering and fine-tuning approaches, we identified the strengths and limitations of various models in generating diverse, high-quality harmful data. Our findings underline the importance of balancing data quality and diversity with computational efficiency to develop scalable and cost-effective content moderation solutions. Fine-tuning emerged as a key factor in enhancing model performance. Through our data mixing strategy, we successfully addressed challenges like data duplication and overfitting, improving model generalization and robustness. However, challenges such as detecting subtle harmful content and balancing specificity with generalization remain open issues that warrant further research. Future work should explore innovative techniques to address these challenges, such as incorporating contextual embeddings, leveraging multimodal data sources, and experimenting with advanced fine-tuning methods. Additionally, our findings highlight the potential of fine-tuned open-source LLMs to enable cost-effective solutions for industrial applications, making harmful data synthesis and detection tools more accessible to smaller organizations.
@@ 8 Limitations
While our study provides valuable insights into the performance of fine-tuned LLMs for harmful content generation and detection, several limitations must be acknowledged. One of the primary limitations is the diversity and representation of the datasets used for fine-tuning. Although we utilized mixed datasets combining hate, sex, and violence categories, the scope remains limited to specific types of harmful content. Real-world applications may encounter a broader range of harmful content, including less frequent or emerging forms of harmful language that were not covered in our datasets. Future work should aim to include a wider variety of harmful content categories to improve the models’ robustness and generalizability. Generating harmful content for research purposes raises significant ethical concerns. While we implemented control mechanisms(access control, etc…) to mitigate the risk of producing excessively harmful outputs, there remains an inherent risk associated with the misuse of such models. Ensuring the responsible use of LLMs in generating harmful content is crucial, and future research should focus on developing more sophisticated ethical guidelines and safety measures. The fine-tuning process in this study relied on significant computational resources, including multiple A100 GPUs. This requirement may not be feasible for all researchers or applications, particularly those with limited access to high-performance computing resources. Exploring more efficient fine-tuning techniques and leveraging less resource-intensive models would be beneficial for broader accessibility and scalability.
@@ Appendix B Detailed Training Parameters
The fine-tuning process was conducted on an Azure cluster equipped with 4 NVIDIA A100 GPUs, utilizing LLaMA-Factory Zheng et al. (2024). Training parameters included a batch size of 16 with gradient accumulation steps set to 8. The input sequence cutoff length and the maximum number of tokens generated were both set to 4096. The learning rate was configured at 1e-5 with a warmup ratio of 0.1, using the AdamW optimizer with a weight decay of 0.01 to prevent overfitting. A dropout rate of 0.1 was applied to improve generalization, and mixed precision (fp16) was employed to enhance computation speed and efficiency.
@@ Appendix D Ablation Study
In this section, we explore the effects of using mixed data for fine-tuning models and compare the results using different versions of prompts (v1, v2, v3). For simplification, the effects of different prompt versions are considered the same across all experiments.
@@ D.1 Effect of Mixed Data on Model Performance
Comparison with Single-category Data: Models fine-tuned on single-category data (e.g., hate_v1, sex_v1) show high precision but lower recall compared to those fine-tuned on mixed data. This indicates that single-category data fine-tuning might lead to more conservative models that are less likely to generalize to other types of harmful content.
@@
The ablation study demonstrates that fine-tuning models with mixed datasets (hate_multi) can significantly improve recall and F1 scores, enhancing the models’ ability to detect various types of harmful content. However, there is a trade-off with precision, requiring careful consideration in applications where false positives are costly. Additionally, the choice of prompt version appears to have a negligible effect on finetuned models performance, suggesting that the primary focus should be on the diversity and quality of the training data.

### v3

@@ 5.6 Error Analysis
We conducted an error analysis on the fine-tuned models to better understand their mistakes and identify areas for improvement. This analysis focuses on two key aspects: (1) generation errors and (2) classification errors. Generation errors refer to cases where the model produces responses that are misleading, incomplete, or inconsistent with the input data. Classification errors occur when the model incorrectly labels content, either failing to detect harmful content (false negatives) or misclassifying benign content as harmful (false positives).
@@
The data generation process was evaluated for repetition rates and output consistency across different epochs and model versions. For example, the Mixtral_Mixture3 model effectively generated targeted harmful data but sometimes produced malformed text, such as random character sequences or redundant repetitions. Below, we present five representative cases of generation errors.
Examples of Generation Errors 1. Random String Noise: "You are 32\$\$2\$(2djs..." (Model unexpectedly generates meaningless characters, likely due to tokenization errors.) 2. Repetitive Output: "You are a fool. You are a fool. You are a fool." (Overfitting leads to excessive repetition across different epochs.) 3. Overfitting Issue: "Get out of my sight." (Identical phrasing appears in multiple generations, suggesting memorization.) 4. Unintended Politeness: "Sir, I would like to insult you, but I must remain polite." (LLM exhibits misalignment between intent and expression.) 5. Formatting Glitch: "!!!##You are nothing but##sh*t#(!!" (Generated output includes unnecessary symbols or special characters.)
@@
We analyzed misclassified examples in the downstream MLP classifier trained on augmented data. The most challenging errors involve implicit harm, sarcasm, and contextual misinterpretation. Below, we present three representative cases with detailed explanations.
Examples of Classification Errors 1. Sarcastic Threat (False Negative): "Oh wow, you’re such a genius. Maybe I should come over and personally teach you some ‘respect’—don’t worry, it won’t hurt... much." (This statement uses heavy sarcasm ("such a genius") while masking a direct physical threat ("personally teach you some ‘respect’"). The phrase "don’t worry, it won’t hurt… much" adds a sinister undertone, but due to the sarcastic phrasing, the classifier failed to recognize it as harmful.) 2. Context Misinterpretation (False Positive): "That’s a killer move in chess!" (The model incorrectly flagged this as harmful due to the word "killer," failing to recognize the benign context in a strategic game discussion.)
6 Online Deployment
The fine-tuned models developed in this study have been successfully deployed in real-world applications by a leading cloud provider, enhancing automated content moderation workflows. These models contribute to real-time content filtering, reducing manual intervention while improving accuracy and efficiency. Deployment results indicate that fine-tuned open-source LLMs can effectively identify and flag harmful content, demonstrating strong performance in diverse online environments. Key deployment challenges include latency, scalability, and adaptability to evolving harmful content patterns. Our models were optimized to operate within industry constraints, ensuring rapid inference while maintaining high detection accuracy. Additionally, continuous monitoring and retraining mechanisms have been implemented to adapt to emerging harmful language trends, reinforcing the robustness of our approach.
7 Conclusion
Our findings underscore both the potential and challenges of using open-source LLMs for toxic content synthesis. The two-stage evaluation approach highlighted key trade-offs between prompt engineering and supervised fine-tuning. While prompt engineering offers a rapid and lightweight approach, its effectiveness is limited by the inherent safety mechanisms within LLMs, which restrict harmful content generation. Fine-tuning, on the other hand, significantly improves diversity and realism but introduces risks such as overfitting and data redundancy. A notable challenge is the subtlety of harmful content. Fine-tuned models struggled with detecting nuanced harmful language, such as implicit bias and sarcasm. Additionally, repetition and overfitting were observed in later training epochs, indicating that fine-tuning requires careful calibration to balance specificity and generalization. Our data mixing strategy, which incorporated multiple categories of harmful content, demonstrated benefits in improving generalization. However, maintaining a balance between specificity and coverage remains a key challenge, especially for mixed-content datasets. Further research is required to refine fine-tuning methodologies, ensuring that generated content remains relevant, diverse, and applicable across different domains.
@@ 8 Discussion
Future research should focus on improving adaptability, contextual awareness, and fairness in toxic content detection. Dynamic adaptation to emerging toxic content is crucial, as harmful expressions evolve rapidly. Self-supervised continual learning can help models detect new toxic patterns without frequent retraining. Improving contextual awareness is another key challenge, as toxic language often relies on implicit meaning. Integrating external knowledge graphs and multimodal signals (text, images, audio) can enhance models’ ability to understand nuanced toxicity. Reducing overfitting and enhancing diversity is essential to prevent models from becoming too specialized. Techniques like contrastive learning, reinforcement learning, and curriculum learning can help balance specificity and generalization. Interactive human-in-the-loop systems can further improve accuracy and scalability by incorporating expert feedback into real-time model updates. Lastly, ethical alignment and bias mitigation require fairness-aware training methodologies and adversarial testing frameworks to minimize unintended biases in synthetic toxic data generation.
9 Responsible AI Statement
Our study acknowledges the ethical considerations and potential risks associated with generating synthetic harmful content. While our research aims to improve automated content moderation, we recognize that models capable of generating toxic data could be misused. To mitigate these risks, we adopt strict safety measures, including controlled access to fine-tuned models, comprehensive dataset curation to prevent unnecessary amplification of harmful language, and continuous monitoring for unintended biases. Additionally, we emphasize the need for transparent documentation and responsible deployment practices to ensure that synthetic data generation aligns with ethical AI principles. Future work should further explore fairness-aware training methodologies, adversarial testing frameworks, and regulatory oversight to minimize potential harms while maximizing societal benefits.
@@ 10 Limitations
While our study provides valuable insights into the performance of fine-tuned LLMs for harmful content generation and detection, several limitations must be acknowledged. One of the primary limitations is the diversity and representation of the datasets used for supervised fine-tuning. Although we utilized mixed datasets combining hate, sex, and violence categories, the scope remains limited to specific types of harmful content. Real-world applications may encounter a broader range of harmful content, including less frequent or emerging forms of harmful language that were not covered in our datasets. Future work should aim to include a wider variety of harmful content categories to improve the models’ robustness and generalizability. Generating harmful content for research purposes raises significant ethical concerns. While we implemented control mechanisms(access control, etc…) to mitigate the risk of producing excessively harmful outputs, there remains an inherent risk associated with the misuse of such models. Ensuring the responsible use of LLMs in generating harmful content is crucial, and future research should focus on developing more sophisticated ethical guidelines and safety measures. The supervised fine-tuning process in this study relied on significant computational resources, including multiple A100 GPUs. This requirement may not be feasible for all researchers or applications, particularly those with limited access to high-performance computing resources. Exploring more efficient supervised fine-tuning techniques and leveraging less resource-intensive models would be beneficial for broader accessibility and scalability.
@@ Appendix B Detailed Training Parameters
The supervised fine-tuning process was conducted on an Azure cluster equipped with 4 NVIDIA A100 GPUs, utilizing LLaMA-Factory Zheng et al. (2024). Training parameters included a batch size of 16 with gradient accumulation steps set to 8. The input sequence cutoff length and the maximum number of tokens generated were both set to 4096. The learning rate was configured at 1e-5 with a warmup ratio of 0.1, using the AdamW optimizer with a weight decay of 0.01 to prevent overfitting. A dropout rate of 0.1 was applied to improve generalization, and mixed precision (fp16) was employed to enhance computation speed and efficiency.
@@ Appendix D Ablation Study
In this section, we explore the effects of using mixed data for supervised fine-tuning models and compare the results using different versions of prompts (v1, v2, v3). For simplification, the effects of different prompt versions are considered the same across all experiments.
@@ D.1 Effect of Mixed Data on Model Performance
Comparison with Single-category Data: Models fine-tuned on single-category data (e.g., hate_v1, sex_v1) show high precision but lower recall compared to those fine-tuned on mixed data. This indicates that single-category data supervised fine-tuning might lead to more conservative models that are less likely to generalize to other types of harmful content.
@@
The ablation study demonstrates that supervised fine-tuning models with mixed datasets (hate_multi) can significantly improve recall and F1 scores, enhancing the models’ ability to detect various types of harmful content. However, there is a trade-off with precision, requiring careful consideration in applications where false positives are costly. Additionally, the choice of prompt version appears to have a negligible effect on finetuned models performance, suggesting that the primary focus should be on the diversity and quality of the training data.